jaythomas:

Evaluate each expression if x = 14, y = 5, and z = 6. 645-01-03-00-00_files/i0070000.jpg a. 645-01-03-00-00_files/i0070001.jpg c. 645-01-03-00-00_files/i0070002.jpg b. 645-01-03-00-00_files/i0070003.jpg d. 645-01-03-00-00_files/i0070004.jpg

justjm:

Can you post the images by going to 'attach file' Or you could drag the image to the textbox Otherwise, I'll assume you're just plugging those values into some expression. Use the order of operations (PEMDAS) to help you through Parenthesis Exponents Multiplication/Division - whichever comes first, either division or multiplication, you do Addition/Subtraction - just like Multiplication/Division...whichever comes first..you do that

CrazyHotGhost:

What even?

jaythomas:

jaythomas:

not working

jaythomas:

nvm thanks tho

CrazyHotGhost:

Np, ig

justjm:

You could also take a screenshot and post it here

jaythomas:

Use the Distributive Property to write each expression as an equivalent expression. Then evaluate the expression. (7 + 6)2 a. 2 × (7 + 6) × 2 = 52 c. 2 × 7 + 2 × 6 = 26 b. 2 × 7 + 2 × 6 = 2 d. 2 × 7 + 6 = 20

justjm:

Do what is inside of the parenthesis first, and then continue on.

CrazyHotGhost:

What even tho?